OpenStudy (anonymous):

Which expression is equivalent to this one? –5(6 + 11)

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 What are the rest of your options? @ihateAtt0t

OpenStudy (anonymous):

A. (–5) • 6 + 11 B. (–5) + 6 + 11 C. (–5) • 6 • 11 D. (–5) • 6 + (–5) • 11

OpenStudy (anonymous):

-5(17) or (-30-55) = -85

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

I just need help answering the problems with the exact answer

TheSmartOne (thesmartone):

Do you know of \(\sf a(b+c)=ab+ac\)

OpenStudy (anonymous):

Sorta

OpenStudy (anonymous):

that is distributive property

OpenStudy (anonymous):

I know what it is but I just need an answer. I will work though. Will give medals.

OpenStudy (anonymous):

the answer is D

TheSmartOne (thesmartone):

So if we use that we get \(\sf a(b+c)=ab+ac\) \(\sf -5(6+11)=(-5\times 6) +(-5\times 11)\) We could also do it by first adding 6+11/ What is \(\sf 6+11=?\)

OpenStudy (anonymous):

yes =17 befor * by -5

OpenStudy (anonymous):

Can you help me with some more questions?

TheSmartOne (thesmartone):

Sure :)

OpenStudy (anonymous):

2(3-1)?

OpenStudy (anonymous):

cant see it

OpenStudy (anonymous):

be fast. have a low Barry

OpenStudy (anonymous):

cant seee the questions

OpenStudy (anonymous):

OMG!!! This is irritating

TheSmartOne (thesmartone):

Refresh if you can't see it.